OpenStudy (anonymous):

What is the area of a square whose perimeter is 100 cm?

OpenStudy (anonymous):

If the area of the square is 100 cm2, then all sides, being of equal length, are the root of 100 = 10. Add the four sides to get the perimeter, 4 x 10 = 40

OpenStudy (anonymous):

perimeter = 4a area of the square = a^2 use this to solve your sum

Parth (parthkohli):

First, solve for one side.\[s + s + s + s = 100\]\[\implies 4s = 100\]\[\implies s = {100 \over 4}\]\[\implies s = 25\]

Parth (parthkohli):

Then find \(s \times s\) which is \(25 \times 25\).

Parth (parthkohli):

@JeremyW24 did it the other way round.

OpenStudy (anonymous):

Did that help? Or did I confuse you more ?

OpenStudy (anonymous):

Thank you, everyone. :)

OpenStudy (anonymous):

No problem @BatmanBane!! Hope that helped you!!